OpenStudy (anonymous):

The perimeter of a rectangle with length, l and width, w is given by P = 2w + 2l. If the perimeter of a certain rectangle is 105 feet and the length is 35 feet, what is the width? Please show all of your work.

sam (.sam.):

P = 2w + 2l 105 = 2w + 2(35) 105 = 2w + 70 35 = 2w 17.5 = w

OpenStudy (anonymous):

thanks for the help i get it now